5^x = 125 (find the exact answer and round 3 decimal points)
5^x = 15 (find the exact answer and round 3 decimal points)

Answer by Alan3354(69443)

5^x = 125 (find the exact answer and round 3 decimal points)
3 (from memory)
-----------
5^x = 15 (find the exact answer and round 3 decimal points)
Then log(5)*x = log(15)
x = [log(15)]/log(5)
x = 1.682606...
x = 1.683